Problem 12-17 Net Present Value Analysis; Internal Rate of Return; Simple Rate of Return (LO12-2, LO12- 3, LO12-6] Casey NelsReg 1 Req 2 Req 3 Req3 Regan Req 4B What is the projects net present value? (Round your final answer to the nearest whole doReg 1 Req 2 Req3 Req 3 Req 4A Req 4A Req 4B Req 4B What is the projects internal rate of return? (Round your answer to wholeReal Req 1 Rea Reg 2 Rea Reg 3 Reqan Req 4A ReqAB Reg 4B What is the projects simple rate of return? (Round your answer to 1
Answer #1
Net operating income 400000
Add: Depreciation 700000
Net cash flows 1100000
1
Now Year 1 Year 2 Year 3 Year 4 Year 5
Investment cost -3500000
Net cash flows 1100000 1100000 1100000 1100000 1100000
Total cash flows -3500000 1100000 1100000 1100000 1100000 1100000
PV factor @ 16% 1 0.862 0.743 0.641 0.552 0.476
Present value of cash flows -3500000 948200 817300 705100 607200 523600
Net present value 101400
2
PV factor internal rate of return=3500000/1100000 = 3.182
The PV factor 3.182 for 5 years is closest to 17%
Internal rate of return = 17%
3
Simple rate of return = Net operating income/Investment cost
Simple rate of return = 400000/3500000= 11.4%
4a
Yes, the company would want Casey to pursue this investment as Net Present value is positive
4b
No, Casey would not be inclined to pursue this investment as as his ROI will decrease